OpenStudy (anonymous):

How many solutions does this system of equations have? x-3y=6 9y=3x-18

OpenStudy (karatechopper):

well, first u want to solve for only one of the variables in one of the equations

OpenStudy (anonymous):

it would have infinite number.Correct?

OpenStudy (anonymous):

??

OpenStudy (karatechopper):

srry im busy

OpenStudy (anonymous):

How many solutions does this system of equations have? x-3y=6=x=3(y+2) x-3y-6=0 y=x/3-2 x=3n y=n-2